OpenStudy (anonymous):

@amistre64 Ok so you used the Geometric sequence: \[B_n=BK^n-P*\text{ Geoemtric sum }\] \[B_n=BK^n-P*\frac{k^n-1}{(k-1}\] Shouldnt it be: \[B_n=BK^n-P*\frac{1-k^n}{1-k}\]

OpenStudy (amistre64):

lol, doesnt matter, you are confusing form with mathical logic \[\frac{k^n-1}{k-1}\implies~\frac{-1}{-1}\frac{k^n-1}{k-1}\implies~\frac{1-k^n}{1-k}\]

OpenStudy (anonymous):

ohh ok gotcha

OpenStudy (anonymous):

Thanks

OpenStudy (amistre64):

its just easier in terms of k in the denominator 1 + r - 1 = r

OpenStudy (anonymous):

ya that i saw

OpenStudy (anonymous):

K thanks :)

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

just for fun ... spose P changes by some rule; \(P_{n+1} = P_{n}(1+r)\), we increase the payment by r% each time. B0 = B B1 = Bk - P B2 = Bk^2 - Pk -P(1+r) B3 = Bk^3 - Pk^2 -P(1+r)k - P(1+r)^2 let 1+r = m, and lets rewrite as B0 = B B1 = Bk - Pk^0m^0 B2 = Bk^2 - Pk^1m^0 - Pk^0m^1 B3 = Bk^3 - Pk^2m^0 - Pk^1m^1 - Pk^0m^2 we do have a pattern can we make a rule for the summation? \[S = k^0m^{n-1} + k^1m^{n-2} + k^2m^{n-3}+...+ k^{n-2}m^1 + k^{n-1}m^0\] \[-km^{-1}S = - k^1m^{n-2} - k^2m^{n-3}-...- k^{n-2}m^1 - k^{n-1}m^0 - k^nm^{-1}\] \[S = \frac{m^{n-1} - k^nm^{-1}}{1-km^{-1}}\] \[B_n = Bk^n - P\frac{m^{n-1} - k^nm^{-1}}{1-km^{-1}}\] \[B_n = Bk^n - P\frac{m^{n} - k^n}{m-k}\] when r=0, m=1 if r=.1/12, then what is our inital basis for a payment? :)

OpenStudy (amistre64):

if the payments are truly variable, can they be determined before hand? or do we setup a variable payment schedule like 10, 12, 15, 20, 21, 32, 33, 36 .... if we setup a payment schedule it may be best to just spreadsheet it

OpenStudy (anonymous):

Well if its variable then i dont think its possible to find a general form You are stuck using the following formula : \[PV=\frac{P_1}{1+r}+\frac{P_2}{(1+r)^2}+...+ \frac{P_n}{(1+r)^n}\]

OpenStudy (anonymous):

Anddd love what you did above ;) I guess thats we can consider that the "variation" with payments :P